9.30: After spending $ 9 comma 900 on​ client-development, you have just been offered a big...

9.30:

After spending $ 9 comma 900 on​ client-development, you have just been offered a big production contract by a new client. The contract will add $ 196 comma 000 to your revenues for each of the next five years and it will cost you $ 99 comma 000 per year to make the additional product. You will have to use some existing equipment and buy new equipment as well. The existing equipment is fully​ depreciated, but could be sold for $ 47 comma 000 now. If you use it in the​ project, it will be worthless at the end of the project. You will buy new equipment valued at $ 27 comma 000 and use the​ 5-year MACRS schedule to depreciate it. It will be worthless at the end of the project. Your current production manager earns $ 80 comma 000 per year. Since she is busy with ongoing​ projects, you are planning to hire an assistant at $ 39 comma 000 per year to help with the expansion. You will have to immediately increase your inventory from $ 20 comma 000 to $ 30 comma 000. It will return to $ 20 comma 000 at the end of the project. Your​ company's tax rate is 35 % and your discount rate is 15.5 %. What is the NPV of the​ contract? Note​: Assume that the equipment is put into use in year 1.
Answer #1
Amount of $9900 spent is sunk cost, not relevant to this analysis
a Net Cash Flow at time 0
After tax Opportunity Cost of not selling existing equipment ($30,550) (47000*(1-0.35)
Investment in new equipment ($27,000)
Net increase in inventory=30000-20000 ($10,000)
Net Cash Flow at time 0= ($67,550)
b Annual Cash Flow
Addition to revenue $196,000
Cost of making ($99,000)
Salary of planning assistant ($39,000)
Annual Before tax cash flow $58,000
Annual after tax cash flow =(58000*(1-0.35) $37,700
c Depreciation Tax shield
Year                       1 2                   3 4                   5
A 5 Year MACRS Depreciation Rate 20% 32% 19.20% 11.52% 11.52%
B=A*27000 Annual Depreciation $5,400 $8,640 $5,184 $3,110 $3,110
C=B*35% Depreciation Tax shield $1,080 $2,765 $995 $358 $358
d Terminal Cash Flow:
Release of working Capital $10,000
e NPV (Net Present Value) of the contract
Present Value of Cash Flow=(Cash Flow)/((1+i)^N)
i= discount rate=15.5%=0.155, N=Year of Cash Flow
N Year 0 1 2 3 4 5
a Initial Cash flow ($67,550)
b Annual Cash Flow $37,700 $37,700 $37,700 $37,700 $37,700
c Depreciation tax shield $1,080 $2,765 $995 $358 $358
d Terminal Cash flow $10,000
CF=a+b+c+d Net Cash Flow ($67,550) $38,780 $40,465 $38,695 $38,058 $48,058 SUM
PV=CF/(1.155^N) Present Value of Net Cash Flow at15.5% discount ($67,550) $33,576 $30,333 $25,114 $21,386 $23,381 $66,239
NPV=Sum of PV Net Present Value $66,238.80